OpenStudy (anonymous):

find an equation of the line that passes through the two given points. (1,7) and (-2, 1)

OpenStudy (mimi_x3):

Find the gradient, by using the gradient formula m = y2-y1/x2-x1 Then use (1,7) and m, to substitute it into y = mx + b , then find b.

OpenStudy (anonymous):

i am so confused

OpenStudy (mimi_x3):

How ? Can't you substitute ? m = 1-7/-2-1 =

OpenStudy (anonymous):

i can right that part out fine, but i am having trouble to find the solution.its -6/-1 right? but where does it go into y=mx+b

OpenStudy (anonymous):

b=-6 and m = -1?

OpenStudy (mimi_x3):

Use x and y to substitute into y=mx +b 7 = 2(1) + b 7 = 2 + b b = 5 The gradient is 2

OpenStudy (mimi_x3):

Do you get it ?

OpenStudy (anonymous):

where did you get 7 from? and the negative 1 is used as an absolute value ?

OpenStudy (mimi_x3):

The coordinate, (1,7) <---x=1, y = 7 -1 ? Absolute value ? There's not absolute value in this question.

OpenStudy (mimi_x3):

no*

OpenStudy (anonymous):

(1,7) and (`2,1) is written 1-7/1--2 , i am wondering how we got rid of the negative and still have an equation on 2 given points

OpenStudy (anonymous):

that should be a negative 2

OpenStudy (mimi_x3):

(1,7) (-2,1) m = y2-y1 / x2-x1 1-7 / -2-1 -6/-3 = 2

OpenStudy (precal):

Note we use the word slope in the US. Gradient is another word for slope. I think it is used in Britain.

OpenStudy (mimi_x3):

Slope and gradient is the same here, I don't know about other countries. Well, I live in Australia so it is used in Australia as well xD

OpenStudy (precal):

In the US we do not use the word gradient. I only know about it because I saw a video (some BBC type video) that used it. I have taught math for about 16 years in public schools and we do not use or test slope using gradient.

OpenStudy (mimi_x3):

Okay, I live in Australia so its different, I never knew that people in the US don't use the word gradient . Well, idk if the poster live in the US or not. What do you use then ? If you don't use the gradient formula ? >_<

OpenStudy (precal):

We use the same formula, it is called the slope formula. As you go up to calculus it is rate of change: delta y over delta x